1. –p → q
  2.  (p → –q) ∨ –q
  3.  p → –(p ∨ q)
  4. (p ∧ q) ∨ (–p ∨ q)
  5. [ (p ∧ q) ∨ –p ] ∧ –q
1
Expert's answer
2022-03-08T11:59:04-0500

1)


"\begin{matrix}
\hline
p & q & -p & -p\rightarrow q \\ \hline
0 & 0 & 1 & 0 \\ \hline
0 & 1 & 1 & 1 \\ \hline
1 & 0 & 0 & 1 \\ \hline
1 & 1 & 0 & 1 \\ \hline
\end{matrix}"

2)


"\begin{matrix}
\hline
p & q & -q & p\rightarrow -q & (p\rightarrow-q)\vee-q \\ \hline
0 & 0 & 1 & 1 & 1 \\ \hline
0 & 1 & 0 & 1 & 1 \\ \hline
1 & 0 & 1 & 1 & 1 \\ \hline
1 & 1 & 0 & 0 & 0 \\ \hline
\end{matrix}"

3)

"\begin{matrix}\hline
p & q & p\vee q & -(p\vee q) & p\rightarrow -(p \vee q) \\ \hline
0 & 0 & 0 & 1 & 1 \\ \hline
0 & 1 & 1 & 0 & 1 \\ \hline
1 & 0 & 1 & 0 & 0 \\ \hline
1 & 1 & 1 & 0 & 0 \\ \hline

\end{matrix}"

4)

"\begin{matrix}\hline
p & q & p \land q & -p & -p \vee q & (p \land q)\vee(-p \vee q)) \\ \hline
0 & 0 & 0 & 1 & 1 & 1 \\ \hline
0 & 1 & 0 & 1 & 1 & 1 \\ \hline
1 & 0 & 0 & 0 & 0 & 0 \\ \hline
1 & 1 & 1 & 0 & 1 & 1 \\ \hline
\end{matrix}"

5)


"\begin{matrix}\hline
p & q & p \land q & -p & (p \land q)\vee-p & -q & {[}(p \land q) \vee -p{]} \land -q \\ \hline
0 & 0 & 0 & 1 & 1 & 1 & 1 \\ \hline
0 & 1 & 0 & 1 & 1 & 0 & 0 \\ \hline
1 & 0 & 0 & 0 & 0 & 1 & 0 \\ \hline
1 & 1 & 1 & 0 & 1 & 0 & 0 \\ \hline
\end{matrix}"